OpenStudy (anonymous):

Finding x intercepts of f(x)= 6sinx-x^2?

OpenStudy (anonymous):

I can get to x^2=6sinx but then how do i reduce it down?

OpenStudy (anonymous):

Can you use a calculator?

OpenStudy (anonymous):

One obvious solution is \(x = 0\), but I don't see how you'll get the other ones (if there are any other ones).

OpenStudy (anonymous):

... without using a calculator, I mean.

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

Gah, had to leave computer. i think my teacher wants this long hand, so no calculator.

OpenStudy (anonymous):

you have to find the x and y intercepts. wouldnt that just give you the y ones?

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

x-intercept = when function crosses x-axis when it does, y=0. solve for x

OpenStudy (anonymous):

right, so 0=6sinx-x^2

OpenStudy (anonymous):

yep.. what are the solutions?

OpenStudy (anonymous):

thats the part im unsure of, what do i do with sinx?

OpenStudy (anonymous):

how about x=0?

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

so, x=0 is one of the x-intercept

OpenStudy (anonymous):

when else can this happen?

OpenStudy (anonymous):

No idea. (0,0) is a point though so thats something

OpenStudy (anonymous):

Dont you do that when you're finding the concavity?

OpenStudy (anonymous):

did you study the numerical methods like Bisection method, Newton-Raphson, Secant method?

OpenStudy (anonymous):

uhhh no.

OpenStudy (anonymous):

without that, analytically, I do not see a way to get the second root.

OpenStudy (anonymous):

or replace sine function by exponentials

OpenStudy (anonymous):

Well, If you can show id be grateful. its possible my teacher forgot to mention.

OpenStudy (anonymous):

\[0={6\over2i}(e^{ix}-e^{-ix})-2x^2\]

OpenStudy (anonymous):

you mean the interpolation?

OpenStudy (anonymous):

using the numerical method?

OpenStudy (anonymous):

yes. to find the umm intercepts

OpenStudy (anonymous):

ok, let us say the second root is approximately at "x=3" you correct it by the following: \[\color{red}{\Large\boxed{x_{\rm new}=x_{\rm old}-\frac{f(x)}{f'(x)}}}\]

OpenStudy (anonymous):

\[x_{\rm new}=3-\frac{6\sin(3)-2(3^2)}{6\cos(3)-4(3)}\approx2.044\]

OpenStudy (anonymous):

How would you get 3?

OpenStudy (anonymous):

oh wait.. i took the wrong function... about the "3", ypu can start with any approximation,.. may be "4".. its fine.. you'd slowly reach the correct value, number of steps will change..

OpenStudy (anonymous):

ugh. this sounds complicated. Is this calculator version any easier?

OpenStudy (anonymous):

would you recommend it?

OpenStudy (anonymous):

\[x_{\rm new}=3-\frac{6\sin(3)-(3^2)}{6\cos(3)-2(3)}\approx2.317\\ x_{\rm new}=2.317-\frac{6\sin(2.317)-(2.317)^2}{6\cos(2.317)-2(2.317)}\approx\boxed{2.21}\]

OpenStudy (anonymous):

the second accurate intercept is 2.2014

OpenStudy (anonymous):

This is called Newton-Raphson interpolation scheme.. if you are taking Calc1, you may aswell understand it and use it!

OpenStudy (anonymous):

haha alright. how do you achieve the 2.317?

OpenStudy (anonymous):

use a calculator! :)

OpenStudy (anonymous):

out of curiosity... check if they meant \[f(x)=6\sin(x-x^2)\]

OpenStudy (anonymous):

well 2.2 is exactly what my calculator graphs it as. there is a space between the x and sin...no parenthesis though.

OpenStudy (anonymous):

I assume if it was 6sin(x-x^2) it would have to have them?

OpenStudy (anonymous):

if it was the way I just said, I'd be very simple to evaluate.. \[0=6\sin(x-x^2)\\ x-x^2={n\pi}\qquad\text{n=0,+-1, +-2, +-3,...}\]

OpenStudy (anonymous):

that would make more sense considering i just did a \[2\sqrt(x)-x \] problem

OpenStudy (anonymous):

Facepalm

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

quadratic. yep.

OpenStudy (anonymous):

you are welcome.